OpenStudy (anonymous):

a brush costs $2 more than a comb.The brush and the comb together cost $3.78.Find the cost of each

OpenStudy (anonymous):

b=c+2 b+c=3.75 Solve the system. Can you do the rest?

OpenStudy (anonymous):

b=c+2 b+c=3.75 c + 2 + c = 3.75 2c + 2 = 3.75 - 2 -2.00 2c = 1. 75 2 2 c = .875 or .88 b = c + 2 b = .88 + 2 b = 2.88